OpenStudy (anonymous):

help me lim as x-->1 \[\left(\begin{matrix}\sqrt[4]{2x-1}+\sqrt[5]{x-2} \\ x-1\end{matrix}\right)\]

OpenStudy (anonymous):

Is is combinatorics?

OpenStudy (anonymous):

same as asking for the derivative of the numerator at x =1

OpenStudy (anonymous):

i am assuming that was a fraction

OpenStudy (anonymous):

what is problem? it makes no sense to me :(

OpenStudy (anonymous):

oh lol

OpenStudy (anonymous):

he/she says lim as x-->1

OpenStudy (anonymous):

that is my assumption, except i meant x goes to 1, not zero

OpenStudy (anonymous):

\[\lim_{x\rightarrow 1}\frac{\sqrt[4]{2x-1}+\sqrt[5]{x-2}}{x-1}\]

OpenStudy (anonymous):

this is asking for \[f'(1)\] where \[f(x)=\sqrt[4]{2x-1}+\sqrt[5]{x-2}\]

OpenStudy (anonymous):

Wow, sat that was fast!

OpenStudy (anonymous):

@lazypig you got this?